OpenStudy (anonymous):

Could someone explain how i can simplify this quadratic equation? I understand that you have to plug it into the quadratic formula. Then, you are to solve it. I did and got out 2 plus or minus the sqaure root of 425 all over 2, but I'm not fully clear on how to simplify this. Could someone please help me? I award medals :))) ;)))

OpenStudy (anonymous):

The eqaution s x^2 -21 = 4x\[x^2 -21 = 4x\]

OpenStudy (surry99):

in order to use the quadratic formula you must put it in the from: ax^2 + bx +c = 0